OpenStudy (anonymous):

We know that the first term is 6 and the fifth term is 96. We also know that n = 5. We need to find the common ratio or r. tn = t1rn-1 96 = 6r5-1 96 = 6r^4 r^4 = 16 r = ±2 please help I don't know where the 6 goes and why it becomes 16

OpenStudy (anonymous):

96 = 6r^4

OpenStudy (anonymous):

okay I understand that 96/6 is 16

OpenStudy (anonymous):

so r^4 = 16 or r^4 =2^4

OpenStudy (anonymous):

after the 16, why does it become two?

OpenStudy (anonymous):

16 =2*2*2*2 = 2^4

OpenStudy (anonymous):

It's the fourth root of 16

OpenStudy (anonymous):

OHHHHHHHokay!

OpenStudy (anonymous):

wait no nvm

OpenStudy (anonymous):

r=2 and since its to the fourth power, it can also be negative

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

but why does r=2 that part isn't clicking for me.

OpenStudy (anonymous):

The fourth root of 16=2

OpenStudy (anonymous):

2^4=16

OpenStudy (anonymous):

2*2*2*2 = 2^4 so r^4 =2^4 so we can say (r/2)^4 =1 now if (r/2) is 1 then it will result in one if raised to power of 4 so r/2=1 so r=2

OpenStudy (anonymous):

But ^4 is even so it can be negative 2 as well

OpenStudy (anonymous):

so it's just something that would go into it??

OpenStudy (anonymous):

similarly one can argue r^4 =(-2)^4 so r can be -2

OpenStudy (anonymous):

Yes, it is a variable this means if we plugged in r=2 2^4=16, 16=16 it is true (-2)^4=16, 16=16 is true that means r=2 or -2

OpenStudy (anonymous):

suppose 3^2 = 9 and now if i ask (x)^2 =9 can you guess what x can be

OpenStudy (anonymous):

okay okay okay! i looked at more of my notes and it explained it! thank you!!!

OpenStudy (anonymous):

3 :)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i hope y get it now